OpenStudy (trisarahtops):

Answer Check!! Medal!! Find the area of the region bounded by the curves y = x2 − 1 and y = cos(x). Give your answer correct to 2 decimal places. I got 1.56

OpenStudy (trisarahtops):

@ganeshie8

OpenStudy (trisarahtops):

@freckles

OpenStudy (xguardians):

Are you allowed calculators on this?

OpenStudy (trisarahtops):

yes

OpenStudy (xguardians):

Graph the two slopes on your calculators, then find the point in which they intersect.

OpenStudy (xguardians):

\[\int\limits_{lowlimit}^{intersect} f(x)\] Of the one that is on top, then do the same thing for the bottom one and subract it.

OpenStudy (xguardians):

In this case, you can just do from 0 to the point of intersection, and after you get your solution, just multiply it by two.

OpenStudy (trisarahtops):

why do I multipy by 2? @xGuardians

OpenStudy (trisarahtops):

if my i have \[\int\limits_{1.175}^{0} \cos(x) - (x^2-1) dx = 1.556\]

OpenStudy (trisarahtops):

@xGuardians

OpenStudy (xguardians):

Why did you put 1.175 on the bottom?

OpenStudy (xguardians):

The reason you multiply by two, is that you need both sides of the y-axis, unless they only wanted one side in the equation.

OpenStudy (trisarahtops):

srry I meant to put it on the top. anyway well then 1.556 *times two is 3.112 @xGuardians

OpenStudy (xguardians):

That should be the correct solutions, good job.